OpenStudy (anonymous):

from a group of six persons, how many committees of three can be formed if two of the six people cannont be on the same committee?

jimthompson5910 (jim_thompson5910):

There are 6 people total, so you can form 6/3 = 2 committees of 3 people each Two people can't be on the same committee. Call these two people person A and person B. So let's say person A is on committee # 1, this means Committee 1: A _ _ and you have 2 slots left and you have 6-2 = 4 people to fill those slots So you have 4 C 2 = 6 ways to fill this committee

jimthompson5910 (jim_thompson5910):

You then double this because you could easily have made person B part of committee # 1 (instead of person A)

OpenStudy (anonymous):

ohhh so its 2(2C1)*4C2 ?

jimthompson5910 (jim_thompson5910):

so I guess you could shorten it to (2 C 1)*(4 C 2)

jimthompson5910 (jim_thompson5910):

2 C 1 = 2, so no need for the extra 2 in front

OpenStudy (anonymous):

i think that only makes out ot 12

OpenStudy (anonymous):

There are (6c3)=20 committees possible. There are (4c1)*(2c2)=4 which include both of the 2 people and are subtracted. 20-4=16

OpenStudy (anonymous):

dont i need the extra two in front cuz person A or person B

jimthompson5910 (jim_thompson5910):

Yes there are 20 committees possible But I think that there are (2 C 1)*(1 C 1)*(4 C 1) = 8 ways to have both A and B on the same committee so there are 20 - 8 = 12 ways to have 2 committees where A and B are not on the same committee

OpenStudy (anonymous):

oh LOLLLL wait the answer is 12 LOLLLL

jimthompson5910 (jim_thompson5910):

yeah thought so

jimthompson5910 (jim_thompson5910):

well there you have 2 ways to get there

OpenStudy (anonymous):

haha dude thanks y'all ... god keep think its 16 so im like...

OpenStudy (anonymous):

the way the question is worded, you are forming a single group of 3 at a time, not 2 groups of 3 at a time. that's where the difference is coming from.

jimthompson5910 (jim_thompson5910):

well whatever you form of one group, the leftover people go into the second group so it's the same thing really

OpenStudy (anonymous):

but it doesn't matter if the 2 people are both in the 2nd group, because that group isn't the committee